OpenStudy (hperez777):

Write the equation of the line, in standard form, that has a y-intercept of 2 and is parallel to 2x + y = -5. Include your work in your final answer.

OpenStudy (hperez777):

@jaelenecastro1998

OpenStudy (sooobored):

parallel lines share the same slope the slope-intercept form of a line is y=mx+b where m is the slope and b is the y intercept

OpenStudy (sooobored):

first identify the slope of 2x+y=-5 any ideas?

OpenStudy (hperez777):

-5 ?

OpenStudy (sooobored):

nope

OpenStudy (hperez777):

no wait 2 right ?

OpenStudy (sooobored):

closer, but no

OpenStudy (hperez777):

-2

OpenStudy (sooobored):

ok, what happens when i subtract -2x from both sides? or move the 2x from the left side to the right side

OpenStudy (sooobored):

yes the slope is -2

OpenStudy (sooobored):

so using the information m=-2 and y-intercept = 2 what would the equation of the line be in slope-intercept form?

OpenStudy (hperez777):

y = -2x - 5

OpenStudy (sooobored):

y-intercept is 2

OpenStudy (sooobored):

not -5

OpenStudy (hperez777):

ohhhh

OpenStudy (hperez777):

y = -2x + 2

OpenStudy (sooobored):

good, final step, the question is asking for the equation in standard form (Ax+By=C) what can you do to turn slope-intercept form into standard form?

OpenStudy (hperez777):

add 2x to the right side so its 2x + y = 2

OpenStudy (sooobored):

yup

OpenStudy (hperez777):

thank youuuuu

OpenStudy (sooobored):

no problem